JustALilCat:

Percy said that any real number for k would cause the system of equations to have no solution. Explain the error in Percy’s statement. 6x + 4y = 14, 3x + 2y = k

Krissy525:

The first thing you would do is eliminate one of the variables (x or y). So you have to use the elimination method. 6x+4y=14 3x(-2)+2y(-2)=k(-2). Then you would get 6x+4y=14 and -6x-4y=-2k Then you divide by two and you get -7, so the answer would be k=all real numbers except -7.